61% of owned dogs in the united states are spayed or neutered. round your answers to four decimal places. if 38 owned dogs are randomly selected, find the probability that a. exactly 23 of them are spayed or neutered. b. at most 26 of them are spayed or neutered. c. at least 23 of them are spayed or neutered. d. between 18 and 23 (including 18 and 23) of them are spayed or neutered.
Answer #1

Answers Pageno het , p be porobability that dogs are şpaged neutered. 091 P-0.6 n = 38 Let a be the random variable for numbeP(xz 23). 38 Pageno (52) (0-61) *(0-22)382 X=23 O-5938 d) P(195x<23)= ele (9)(0-6)*(0.39%. X=18 = P(xsa3)-P(x318) = 0.5374-0.